# rQUFguide.bib

### v1

@@ Abstract
We investigate the disparity in the microstructural properties of the Limit Order Book (LOB) across different relative tick sizes. Tick sizes not only influence the granularity of the price formation process but also affect market agents’ behavior. A key contribution of this study is the identification of several stylized facts, which are used to differentiate between large, medium, and small tick stocks, along with clear metrics for their measurement. We provide cross-asset visualizations to illustrate how these attributes vary with relative tick size. Further, we propose a Hawkes Process model that accounts for sparsity, multi-tick level price moves, and the shape of the book in small-tick stocks. Through simulation studies, we demonstrate the universality of the model and identify key variables that determine whether a simulated LOB resembles a large-tick or small-tick stock. Our tests show that stylized facts like sparsity, shape, and relative returns distribution can be smoothly transitioned from a large-tick to a small-tick asset using our model. We test this model’s assumptions, showcase its challenges and propose questions for further directions in this area of research.
@@ 1 Introduction
The rapid electronification of the financial market has made the data structure Limit Order Books the central domain of essentially all trading, particularly for equities. Limit Order Books match sellers with the buyers according to their price priority first, and queue priority second. These markets are classified as order driven instead of quote driven since there exists a lit, or in other words publicly visible, list of unmatched orders that any market participant can utilize for their trading needs. There are largely three kinds of orders: Market Orders (MOs), Limit Orders (LOs) and Cancel Orders (COs). Limit Orders are representative of a market agent willing to buy / sell a certain quantity of the security at a certain price. Cancel Orders are cancellations of existing LOs by the market agent. Market Orders are similar to Limit Orders however their price is aggressive enough to match quotes on the opposite side of the book (for eg. a buy MO will match with the LOs at ask side of the LOB). A lot of variations, provided by the exchange to their customers for a variety of reasons, exist in these three order types however these orders, in their essential components, have an order price and order size. The order price defines the price at which the agent placing the order is willing to buy or sell the asset. The granularity of the order price in modern lit LOBs is set by a quantity known as the tick-size. By definition it is the minimum difference in an order’s price. Tick-sizes vary both by asset classes and markets. Focusing on equities, the tick-sizes are set to be constant in the US equities exchanges uniformly for all stocks [^1] Certain penny stocks have a tick-size of one hundredth of a cent however their market capitalisation is not significant. to one cent (or 0.01\$). However in Europe, the tick-sizes change with the price of each security on a ladder like logic. There are several discussions in the literature on which tick-size regime is optimal for the welfare of market participants. We will focus on US securities in this paper.
@@
We define the ratio of the tick-size to the price (in \$) as the relative tick-size. In this work, we showcase the fact that with the constant tick-size of the US securities and with a wide range of associated prices of the security, the LOB behaves significantly differently with varying relative tick-size. The aim of this discussion is to then highlight the challenges of modelling LOBs which have a very low relative tick-size. We therefore categorize a stock into large and small tick stock where a large tick stock corresponds to a relatively high relative tick-size (of the order of $10^{-4}\$$) and a small tick stock corresponds to a lower relative tick-size (of the order of $10^{-5}\$$ or lower). We tackle the problem of simulating a general LOB, regardless of its relative tick-size, by modelling the LOB dynamics using a point process.
@@
    We demonstrate a set of stylized facts and define clear metrics of measuring the stylized facts.
@@
    We propose a simple extension to the Compound Hawkes Model in [jain2023hawkes] to account for a sparse LOB with a wide spread.
@@
    We showcase the universality of this model (from small to large tick assets) by depicting various attributes of the LOB by varying the parameters of this model.
@@
    We challenge the assumptions of this model and propose future research directions to tackle them within our model’s framework.
The paper is organized as follows. In Section 2, we summarize the current state of the art on LOB modelling, particularly focusing on the Hawkes Process methods of doing so. We also provide a study of the existing literature concerned with small tick stocks. In Section 3, we develop several stylized facts and their metrics. We provide cross-asset visualizations of these facts and their metrics to depict the various classes and their corresponding stylized facts. In Section 4 we extend a Hawkes Process model of the LOB with a few additional state variables and parameters to account for sparsity in the LOB and multi-tick wide spread. In Section 5, we present some simulation results of a set of random parameters for the above mentioned model. We show the fact that our model is flexible enough to account for several stylized facts of all kinds of relative tick-size. We also compare the performance to a Poisson Process model to track the value of using a Hawkes Process model. Finally, in Section 6, we test the assumptions of the model and provide evidence for and counter-evidence against these assumptions. Section 7 concludes and provides a set of open research questions for the LOB modelling research community.
2 Related Work
2.1 Limit Order Book Modelling:
[Gould2013] conducted a comprehensive review of Limit Order Books (LOBs), examining their characteristics and presenting various models for LOB simulation. Similarly, [cont2011] provided a survey that highlighted the effectiveness of several zero-intelligence models in LOB modeling, offering empirical observations to validate the models’ outputs. For a more detailed exploration of the microstructural statistics of the LOB and the associated modeling techniques, the work by [abergel_anane_chakraborti_jedidi_munitoke_2016] serves as a key reference. [jain2024limit] is a more recent review with a breakdown on the basis of the core methodology used, and the stylized facts tested. They note that the field of order book simulations continues to evolve in parallel with advancements in modeling techniques. The emergence of deep learning has led to the development of numerous architectures aimed at replicating the properties and dynamics of the order book. By far, however, the modelling of the LOB has been constrained to that of large tick stocks and there is a significant gap in the field of LOB modelling when we are concerned about small tick stocks.
2.2 Full LOB Models:
We categorize an LOB model as a full LOB model if the modelling method is general enough to be applied to deeper dimensions rather than just the top of the book. We review the mathematical methodology of some of the state of the art below.
[Abergel2015], using the exponential kernels of a multivariate Hawkes Process, model a moving frame of the LOB with $K$ price limits on each side of the LOB measured from the opposite best quote. Therefore the frame moves with events at both sides. They model the MOs and LOs (at the $2K$ levels) to create a $22K+$ 2 dimensional multivariate Hawkes Process. For Cancels, they use a doubly stochastic Poisson Process proportional to the current queue size for a level. They formulate the dynamics of a queue size at a level $i$ for $i\in 1,\cdots,K$ for each side by using the affect of MOs, LOs, COs affecting level $i$ and shift operators for when the events on the other side of the LOB change the number ing of the levels. They develop the generator operator and thereafter long time limits of the price and volume processes from this model.
[cont2021stochastic] form an SPDE model for centred volume density $x$ ticks away from the mid-price $:=u_{t}(x)$. They model the arrival of limit orders by a function of $x$. They represent MOs using the same function i.e. as marketable limit orders. Cancels are handled in three ways: firstly, outright cancels as constant multipliers on the current volume density, secondly, cancel and replace randomly at nearby levels as a diffusion term, and finally, cancel and replace towards mid price as a convection term on the current volume density. Finally they add a multiplicative Wiener process term to the density process for modelling the HFT orders’ submissions and cancels.
[horst2019scaling] develop step functions for instantaneous volume density on each side of the book, as functions of time and an absolute price coordinate $y$. They assume MO size to be exactly equal to the current top of the book’s size and in-spread orders improve the price by one tick only. MOs and IS orders (denoted as active orders) have an intensity as a separable function of a function of the current mid-price and a time dependent intensity function. LOs and COs (denoted as passive orders) have a size $z$ and occur at a distance $x$ away from the top of the book (on the respective side) with some intensity over $t,x$ and a probability distribution function of $z$. Their framework is quite general and they formulate the price and volume dynamics within this framework and study their long time limits.
2.3 Hawkes Process Models:
The Hawkes Process has emerged as a promising model for addressing the limitations of Poisson Processes in modeling Limit Order Book (LOB) queueing systems ([jain2024limit]). [Bacry2015] provide a comprehensive review, outlining the fundamental concepts of the Hawkes Process, including its mathematical foundations, key properties, and various applications, with a particular focus on Order Book models.
Recent studies, such as those by [Hawkes2018], have further explored the financial applications of Hawkes Processes, underscoring their utility in modeling a range of market phenomena. A significant area of discussion in the literature concerns the selection of appropriate kernel functions.
The extension to multidimensional Hawkes Processes has gained traction in LOB modeling, leading to the development of various formulations. [toke2010market] introduces a two-agent model employing one-dimensional Hawkes Processes for Market Orders and Limit Orders. [bacry2016estimation] take a different approach by categorizing order book events and utilizing an eight-dimensional Hawkes Process to model the bid and ask sides separately. [kirchner2017estimation] propose a non-parametric estimation method for Hawkes Processes, optimizing hyperparameters using the AIC statistic. [jain2023hawkes] extend the non-parametric estimation method for slowly decaying kernels. Alternatively, [fonseca2014calib] propose a generalized method of moments for fitting Hawkes Processes, offering a faster parameter estimation compared to traditional maximum likelihood methods. They validate their approach using key stylized facts and compare the results with MLE baselines. [jain2023hawkes] develop a Compound Hawkes Process with spread closing events being state-dependent to model the LOB while being cognizant of the distribution of individual orders’ sizes. They showcase the fact that their methodology is general enough to be applied to all kinds of stocks however due to their model assumptions being too strong for small tick stocks, their model fails in matching the stylized facts of small tick equities.
2.4 Affect of tick-size on LOB models:
As noted in [LaSpada2011], the relation between tick-size and short term microstructural dynamics of a stock’s LOB has been well documented in the literature. They also establish certain facts of the long-term behaviour of the price process as they vary the tick-size. We also refer to the book by [bouchaud2018trades] for a detailed analysis of some stylized facts and their variation against relative tick-size. They note that the tick-size constrains the bid-ask spread to a large extent which also enforces the price changes to be almost purely driven by queue-depletions in large-tick assets. Moreover, they find that the top of the book volumes in large-tick stocks constitute a significant fraction of the daily traded volume however for small-tick stocks, this fraction goes down to the order of $10^{-2}$. They further find that the effect of queue-position and sizes of market orders relative to existing top of the book queue size are remarkably different in the two categories as well. In addition to this, they show that a significant fraction of the daily traded volume in small tick stocks is from hidden orders in the LOB. The number (higher for large tick), individual size (lower for large tick) and frequency (much higher in large tick) of top of the book orders also is remarkably different between large and small tick stocks. In terms of stylized facts tested, they make use of the bid-ask spread (as a function of time-of-day and its empirical distribution), LO arrival rates as a function of distance from the top of the book, order sizes’ distributions, mean shape of the LOB, volume at top of the book compared with mean market order size and daily traded volume and more to differentiate the microstructural properties between large and small tick stocks. Similarly, [briola2024deep], also test the distribution of bid-ask spread, volumes at the top of the book and the distance between the top of the book and the 10th best top of the book against 3 categories of relative tick-size. They find that each of these three stylized facts differ remarkably across these three categories. Further in [briola2024hlob], the authors show that the mutual information of the LOB top 10 best queues on both sides of the book have significantly different structures across these 3 categories of relative tick-size. This is evidence of the requirement of a different modelling approach for small tick assets.
In the following section we perform visualisations a series of stylized facts and try to formulate the exact quantity in which there is a clear trend as we decrease the relative tick-size.
3 Stylized Facts
@@
To assess the variations in microstructural properties of the Limit Order Book (LOB) across different relative tick sizes, we analyze several key aspects of LOB states over a one-year period, and report their empirical characteristics. These characteristics are commonly referred to in the literature as stylized facts, which represent consistent empirical regularities observed in financial markets ([bouchaud2018trades]). The selection of stylized facts is intended to ensure that the most salient features of the LOB microstructure are captured within this statistical framework, thereby facilitating a comprehensive understanding of its dynamic behavior ([cont2011]). We present an analysis of several stylized facts for 15 stocks. This collection (Table 1), to facilitate a cross category analysis, has price varying from 6\$ to 1800\$ and average bid-ask spread varying from 1 tick to 100s of ticks. For each stock, high-resolution, tick-by-tick LOB data obtained from the LOBSTER provider [LOBSTER] is employed. Notably the US equities have a constant tick size of one cent. We define relative tick size as the ratio between the tick size and the price of the security.
@@
We categorize stocks based on their average spread $\overline{s}$ as follows:
@@
Stock Avg Spread (in ticks) Avg Price (in \$) Rel. Tick-Size (in bps) Category
@@
The stylized facts we showcase are as follows.
3.1 Bid-Ask Spread:
Figure 1: Mean Spread(in ticks) vs Relative Tick Size(in bps) : Annotations depict the stock’s mean and variance of the spread. Remarkably, the average spread seems to be highly correlated with the relative tick size, or, as noted by [bouchaud2018trades], by price. There are clearly two regimes here - since the spread can never be less than one tick, there is a flattening of this relation for large tick stocks, followed by a steady but noisy change to an almost linear (in log-log scale) relationship between the two for small and medium tick stocks.
As noted in [jain2024limit], building a realistic and general Limit Order Book (LOB) model requires careful replication of the empirical density of spreads, both around the median and at the tails, as well as capturing the time-of-day dependence of spreads. This applies across stocks with varying relative tick sizes. Additionally, the price of a security plays a significant role in determining its mean spread, with some non-linearity observed in large-tick stocks due to the constraint that the spread cannot fall below one tick, as seen in [bouchaud2018trades] and in Figure 1.
Let the bid-ask spread $s(t)$ be the difference between the best ask price $p_{\text{ask}}(t)$ and the best bid price $p_{\text{bid}}(t)$ at time $t$ Mathematically:
$s(t)=p_{\text{ask}}(t)-p_{\text{bid}}(t)$
We plot the average spread against the average price on a log-log scale to explore the relationship between price and spread.
We observe a gradual increase in the mean bid-ask spread as the relative tick size decreases across stocks in the dataset. For instance, the mean spread for INTC, a large-tick stock, is $1.02$, whereas for AMZN, a small-tick stock, it reaches $47.18$. In the log scale, the tail of the spread distribution is more pronounced for AMZN, with significant values persisting over dozens of ticks, while for INTC, it decays to near zero after just two ticks (Figure 21). To measure the difference in shapes of the spread distributions, we make use of the Weibull distribution as our candidate distribution. This choice is motivated by the empirical fact that with decreasing relative tick size, one observes the mode shifting from 1 tick to multiple ticks however the tail of the distribution dies quicker as well. Considering this, the Weibull distribution is a prime candidate. We make use of the method of moments to calculate the Coefficient of Variation or CV (i.e. variance divided by mean squared) of the empirical observed spreads for one calendar year. CV is governed by the shape parameter of the Weibull distribution alone ([weibull]) and therefore is a good statistic to measure the differences in shape across stocks. We show the relation of CV to the relative tick size in Figure 2 and we fit a linear regression line on this scatter plot (shown in red). One can clearly observe a definitive trend as we decrease the relative tick size, the slope of the red line is $-2.721$, and so we have the relation $CV\propto\text{(Rel. Tick-Size)}^{-2.7}$. Therefore we conclude that the shape of the spread distribution is directly related to the relative tick size of a stock.
@@ 3.1 Bid-Ask Spread:
In the time-of-day analysis, we observe tick-constrained dynamics in large-tick stocks like INTC, where the spread remains floored at one tick throughout the trading day, despite price discovery causing fluctuations in the spread. This behavior is markedly different in medium and small-tick stocks, where the spread exhibits more variation throughout the day.
3.2 Price Changes:
Figure 3: Mean of Trade induced Price Moves (in \$): As we move from large tick stocks to medium and small tick, we exit the domain of half tick mid price moves to multiple ticks. This is an important observation since, for example, the volatility process (for absolute returns) is scaled by the mid-price changes’ scale which, as we show here, is determined by the rel. tick size. The slope of the small/medium tick fit is approximately $-1$. We therefore conclude that price moves are scaled by the current price which is a tenet in Geometric Brownian Motion models of the price process.

### v2

@@ Abstract
Tick sizes not only influence the granularity of the price formation process but also affect market agents’ behavior. We investigate the disparity in the microstructural properties of the Limit Order Book (LOB) across different relative tick sizes. A key contribution of this study is the identification of several stylized facts, which are used to differentiate between large, medium, and small tick stocks, along with clear metrics for their measurement. We provide cross-asset visualizations to illustrate how these attributes vary with relative tick size. Further, we propose a Hawkes Process model that accounts for sparsity, multi-tick level price moves, and the shape of the book in small-tick stocks. Through simulation studies, we demonstrate the universality of the model and identify key variables that determine whether a simulated LOB resembles a large-tick or small-tick stock. Our tests show that stylized facts like sparsity, shape, and relative returns distribution can be smoothly transitioned from a large-tick to a small-tick asset using our model. We test this model’s assumptions, showcase its challenges and propose questions for further directions in this area of research.
@@ 1 Introduction
The rapid electronification of the financial market has made the data structure Limit Order Books the central domain of essentially all trading, particularly for equities. Limit Order Books match sellers with the buyers according to their price priority first, and queue priority second. These markets are classified as order driven instead of quote driven since there exists a lit, or in other words publicly visible, list of unmatched orders that any market participant can utilize for their trading needs. There are largely three kinds of orders: Market Orders (MOs), Limit Orders (LOs) and Cancel Orders (COs). Limit Orders are representative of a market agent willing to buy / sell a certain quantity of the security at a certain price. Cancel Orders are cancellations of existing LOs by the market agent. Market Orders are similar to Limit Orders however their price is aggressive enough to match quotes on the opposite side of the book (for eg. a buy MO will match with the LOs at ask side of the LOB). A lot of variations, provided by the exchange to their customers for a variety of reasons, exist in these three order types however these orders, in their essential components, have an order price and order size. The order price defines the price at which the agent placing the order is willing to buy or sell the asset. The granularity of the order price in modern lit LOBs is set by a quantity known as the tick-size. By definition it is the minimum difference in an order’s price. Tick-sizes vary both by asset classes and markets. Focusing on equities, the tick-sizes are set to be constant in the US equities exchanges uniformly for all stocks [^1] Certain penny stocks have a tick-size of one hundredth of a cent however their market capitalisation is not significant. to one cent (or 0.01\$). However in Europe, the tick-sizes change with the price of each security on a ladder like logic. We define the ratio of the tick-size (in \$) to the price (in \$) as the relative tick-size. We report this ratio in units of bps which is equal to $10^{-4}$. There are several discussions in the literature on which tick-size regime is optimal for the welfare of market participants. We will focus on US securities in this paper.
1.1 Affect of tick-size on LOB dynamics:
As noted in [LaSpada2011], the relation between tick-size and short term microstructural dynamics of a stock’s LOB has been well documented in the literature. They also establish certain facts of the long-term behaviour of the price process as they vary the tick-size. We also refer to the book by [bouchaud2018trades] for a detailed analysis of some stylized facts and their variation against relative tick-size. To differentiate between large- and small-tick stocks, they analyze various microstructural characteristics, including the bid-ask spread, limit order arrival rates, order size distributions, LOB shape, and top-of-book volumes relative to daily traded volumes. Similarly, [briola2024deep], also test the distribution of bid-ask spread, volumes at the top of the book and the distance between the top of the book and the 10th best top of the book against 3 categories of relative tick-size. They find that each of these three stylized facts differ remarkably across these three categories. Further in [briola2024hlob], the authors show that the mutual information of the LOB top 10 best queues on both sides of the book have significantly different structures across these 3 categories of relative tick-size. This is evidence of the requirement of a different modelling approach for small tick assets.
@@ 1.1 Affect of tick-size on LOB dynamics:
In Section 2 we perform visualisations for a series of stylized facts and try to formulate the exact quantity in which there is a clear trend as we decrease the relative tick-size. To the best of our knowledge, this is the first work which defines quantifiable statistics of the LOB dynamics and depict their variation with rel. tick-size of several assets. In this work, we showcase the fact that with the constant tick-size of the US securities and with a wide range of associated prices of the security, the LOB behaves significantly differently with varying relative tick-size. The aim of this discussion is to then highlight the challenges of modelling LOBs which have a very low relative tick-size. We therefore categorize a stock into large and small tick stock where a large tick stock corresponds to a relatively high relative tick-size (of the order of a few bps) and a small tick stock corresponds to a lower relative tick-size (of the order of 0.1 bps or lower).
1.2 Limit Order Book Modelling:
[Gould2013] conducted a comprehensive review of Limit Order Books (LOBs), examining their characteristics and presenting various models for LOB simulation. Similarly, [cont2011] provided a survey that highlighted the effectiveness of several zero-intelligence models in LOB modeling, offering empirical observations to validate the models’ outputs. For a more detailed exploration of the microstructural statistics of the LOB and the associated modeling techniques, the work by [abergel_anane_chakraborti_jedidi_munitoke_2016] serves as a key reference. [jain2024limit] is a more recent review with a breakdown on the basis of the core methodology used, and the stylized facts tested.
Particularly, they note that the Hawkes Process has emerged as a promising model for addressing the limitations of Poisson Processes in modeling Limit Order Book (LOB) queueing systems. [Bacry2015] provide a comprehensive review, outlining the fundamental concepts of the Hawkes Process, including its mathematical foundations, key properties, and various applications, with a particular focus on Order Book models. Recent studies, such as those by [Hawkes2018], have further explored the financial applications of Hawkes Processes, underscoring their utility in modeling a range of market phenomena. The extension to multidimensional Hawkes Processes has gained traction in LOB modeling, leading to the development of various formulations. [toke2010market] introduces a two-agent model employing one-dimensional Hawkes Processes for Market Orders and Limit Orders. [bacry2016estimation] take a different approach by categorizing order book events and utilizing an eight-dimensional Hawkes Process to model the bid and ask sides separately. [kirchner2017estimation] propose a non-parametric estimation method for Hawkes Processes, optimizing hyperparameters using the AIC statistic. [jain2023hawkes] extend the non-parametric estimation method for slowly decaying kernels. They develop a Compound Hawkes Process with spread closing events being state-dependent to model the LOB while being cognizant of the distribution of individual orders’ sizes. They showcase the fact that their methodology is general enough to be applied to all kinds of stocks however due to their model assumptions being too strong for small tick stocks, their model fails in matching the stylized facts of small tick equities.
By far the modelling of the LOB has been limited to that of large tick stocks and there is a significant gap in the field of LOB modelling when we are concerned about small tick stocks. We categorize an LOB model as a full LOB model if the modelling method is general enough to be applied to deeper dimensions rather than just the top of the book. We review the mathematical methodology of some of the state of the art below.
[Abergel2015], using the exponential kernels of a multivariate Hawkes Process, model a moving frame of the LOB with $K$ price limits on each side of the LOB measured from the opposite best quote. Therefore the frame moves with events at both sides. They develop the generator operator and thereafter long time limits of the price and volume processes from this model. Using a different approach, [cont2021stochastic] propose an SPDE (Stochastic Partial Differential Equations) model for the centered volume density $u(x,t)$, representing volume at $x$ ticks from the mid-price. Limit orders and market orders are modeled as functions of $x$, with cancellations handled in three ways: constant multipliers over Limit Order intensities, diffusion for cancel-and-replace near existing levels, and convection towards the mid-price for liquidity seekers. A multiplicative Wiener process is added to capture high-frequency trading activity. Similarly, [horst2019scaling] develop step functions for instantaneous volume density on each side of the book, as functions of time and an absolute price coordinate $y$. Their framework is quite general and they formulate the price and volume dynamics within this framework and study their long time limits.
While these SPDE models are useful for calculations of the steady state behaviour of a general LOB, we still stress the gap in literature on simulating a discrete price grid general LOB which is faithful to the stylized facts of the LOB. We tackle the problem of simulating a general LOB, regardless of its relative tick-size, by modelling the LOB dynamics using a point process.
@@ 1.2 Limit Order Book Modelling:
    In Section 2, we demonstrate a set of stylized facts and define clear metrics of measuring the stylized facts.
@@
    We propose a simple extension in Section 3 to the Compound Hawkes Model in [jain2023hawkes] to account for a sparse LOB with a wide spread.
@@
    We showcase the universality of this model (from small to large tick assets) by depicting various attributes of the LOB by varying the parameters of this model in Section 4.
@@
    Finally, in Section 5, We challenge the assumptions of this model and propose future research directions to tackle them within our model’s framework.
2 Stylized Facts related to Tick-Size
@@
To assess the variations in microstructural properties of the Limit Order Book (LOB) across different relative tick sizes, we analyze several key aspects of LOB states over a one-year period, and report their empirical characteristics. These characteristics are commonly referred to in the literature as stylized facts, which represent consistent empirical regularities observed in financial markets ([bouchaud2018trades]). The selection of stylized facts is intended to ensure that the most salient features of the LOB microstructure are captured within this statistical framework, thereby facilitating a comprehensive understanding of its dynamic behavior. We present an analysis of several stylized facts for 15 stocks over a whole year’s LOB data (2nd January 2019 to 30th December 2019) i.e. 250 trading days. This collection (details in Table 1), to facilitate a cross category analysis, has average mid-price varying from 6\$ to 1800\$ and average bid-ask spread varying from one tick to hundreds of ticks. Notably the US equities have a constant tick size of one cent [^2] Some stocks have a tick size of one-hundredth of a cent however their market capitalization is very small compared to the stocks considered in this study.. For each stock, high-resolution, tick-by-tick LOB data obtained from the LOBSTER provider [LOBSTER] is employed. The stylized facts we showcase are as follows. Most of the statistics below will be studied with respect to the ‘relative’ tick-size $\equiv\epsilon$ which is defined as the ratio between the tick size and the average mid-price of the security. We choose relative tick-size as our measure for this study rather than average mid-price (as done in [bouchaud2018trades]) since we posit that it is the ratio of tick-size to the price of the security which is the driver of microstructural disparities and not the price alone. Therefore, our analysis would be applicable across other exchanges where tick-sizes are not constant uniformly. One such set of exchanges are the stock exchanges in Europe (London Stock Exchange, Xetra Frankfurt, etc. have a ladder-like logic for a dynamic tick-size). We report $\epsilon$ in bps (i.e. one hundredth of a percent) like in the below equation. Here $<P_{mid}>$ (in \$) is the average mid price of the security over the one year period. Denoting tick-size by $\delta$ (which is a constant and equal to one cent for this work), we have:
$\epsilon$ $:=\frac{\delta}{<P_{mid}>}\times 10^{4}\text{bps}$ (1)
2.1 Bid-Ask Spread : Mean and Coefficient of Variation vs Relative Tick-Size
Bid-Ask Spread (or simply ‘spread’) of an asset is determined by the difference between the best offer and the best bid. It thus represents the premium one has to pay if they desire to ascertain their order’s fill instead of posting a passive order. This leads to the intuitive expectation that the spread should scale with the price, as traders typically perceive the premium in terms of a proportion of the security’s value. It is therefore natural to examine the variation of spread with respect to the price of the security as done in [bouchaud2018trades]. In this study, in order to motivate the affect of tick-sizes on the LOB dynamics, we focus on its variation with relative tick-size which itself is proportional to inverse of the price.
Let the bid-ask spread $s(t)$ be the difference between the best ask price $p_{\text{ask}}(t)$ and the best bid price $p_{\text{bid}}(t)$ at time $t$ We measure $s(t)$ in ticks rather than dollars to stress on the number of price levels spanned by the bid-ask spread of the stock. Mathematically:
$s(t)=\frac{p_{\text{ask}}(t)-p_{\text{bid}}(t)}{\delta}$ (2)
The spread process is a discrete time random process and so for $i=1,...,N$, at a time $0\leq t_{i}\leq T$, with $T$ being the end-of-day time in seconds for the trading day, the spread process remains constant with value $s(t_{i})$ till the next time $t_{i+1}$. It is therefore necessary to calculate any statistic over this discrete time random process in a time-weighted manner. We define the time-weighted average (TWA) spread (denoted by $\overline{s}$) as follows. Moving forwards, we distinguish a time-weighted average statistic from a uniform-weighted average statistic by the notation of $\overline{a}$ for TWA over quantity $a$. The uniformly-weighted average statistics will be reported using the notation $<a>$.
$\overline{s}:=\frac{\sum^{N}_{i=1}s(t_{i})\Delta t_{i}}{T}$ (3)
@@ 2.1 Bid-Ask Spread : Mean and Coefficient of Variation vs Relative Tick-Size
We categorize the stocks presented in Table 1 based on their average spread $\overline{s}$ as follows:
@@
Stock Avg Spread (in ticks) Avg Mid-Price (in \$) Rel. Tick-Size (in bps) Category
@@
Figure 1: Average Spread (in ticks) vs Relative Tick Size(in bps) : Remarkably, the average spread seems to be highly correlated with the relative tick size, or, as noted by [bouchaud2018trades], by mid-price. There are clearly two regimes here - since the spread can never be less than one tick, there is a flattening of this relation for large tick stocks, followed by a steady but noisy change to an almost linear (in log-log scale) relationship between the two for small and medium tick stocks. The slope of the orange fit is approximately $-1$. Therefore we can conclude that the Average Spread is proportional to the Average Mid-Price.
In Figure 1, we plot the average spread in ticks against the average price on a log-log scale to explore the relationship between price and spread. We observe a gradual increase in the mean bid-ask spread as the relative tick size decreases across stocks in the dataset. For instance, the mean spread for INTC, a large-tick stock, is $1.02$ ticks, whereas for AMZN, a small-tick stock, it reaches $47.18$ ticks. We find two regimes here, for large tick stocks, the spread is almost always constant and equal to one tick (hence they are also called tick-constrained stocks). However for medium and small tick stocks, there seems to be a power law between $\overline{s}$ and $\epsilon$ with an exponent of $-1.23$ in our dataset. Thus, we state the following relation approximately. We note that a similar relation was found in [bouchaud2018trades].
$\overline{s}$ $\propto\epsilon^{-1}$ (4)
@@
Focusing now on the instantaneous values of the spread-in-ticks ($s(t)$), we plot the probability distribution function in the Appendix A (Figure 20). We note that the distribution function varies significantly among the three categories of stocks in both scale and shape. To measure the difference in shapes of the spread distributions, we make use of the Weibull distribution as our candidate prior. This choice is motivated by the empirical fact that with decreasing relative tick size, one observes the mode shifting from one tick to multiple ticks however the tail of the distribution dies quicker as well. Considering this and the fact that this distribution is both simple and versatile, the Weibull distribution is a suitable candidate. We make use of the method of moments to calculate the Coefficient of Variation or CV (i.e. variance divided by mean squared) of the empirical observed spreads for one calendar year. CV is governed by the shape parameter of the Weibull distribution alone ([weibull]) and therefore is a good statistic to measure the differences in shape across stocks. We show the relation of CV to the relative tick size in Figure 2 and we fit a linear regression line on this scatter plot (shown in red). One can clearly observe a definitive trend as we decrease the relative tick size, the slope of the red line is $-2.721$, and so we have the relation $CV\propto\text{(Rel. Tick-Size)}^{-2.7}$. Therefore we conclude that the shape of the spread distribution is directly related to the relative tick size of a stock. Note that unlike the previous figure, there does not seem to be a regime change when we move from large to small tick stocks. One potential reason for this could be that both $\sigma^{2}$ and $\mu^{2}$ have two regimes due to the minimum value of spreads in large-tick stocks and when we take the ratio, these two slopes cancel each other out giving us a single smooth regression against $\epsilon$.
2.2 Mid-Price Changes: Probability Distribution vs Relative Tick-Size
Mid-Price changes can occur either due to a new best bid or ask order; or due to a market order or cancellation depleting the top of the book volume at bid or ask. One intuitively expects the market participants to improve the price one tick at a time (since they pay one tick of premium to improve their fill probability). Additionally, a queue depletion would mean that the next best quote’s price level would be one tick away from the previous best. Therefore, intuitively, mid-price changes should happen, mostly, at half-tick increments. We show in the following that while this is true for large tick stocks, the average mid-price changes scales with the inverse of the relative tick size for the rest of the stocks we study.
Figure 3: Mean of Trade induced Mid-Price Changes (in \$): As we move from large tick stocks to medium and small tick, we exit the domain of half tick mid price moves to multiple ticks. This is an important observation since, for example, the volatility process (for absolute returns) is scaled by the mid-price changes’ scale which, as we show here, is determined by the rel. tick size. The slope of the small/medium tick fit is approximately $-1$. We therefore conclude that price moves are scaled by the current price which is a tenet in Geometric Brownian Motion models of the price process.
To measure the behavior of mid-price changes in the Limit Order Book (LOB), we analyze the impact of trades on the LOB. We show the mean trade induced price moves (denoted by $<r_{mid}>:=\mathbb{E}_{(trades)}\Delta p_{mid}(t)$, where $p_{mid}(t)$ is the instantaneous mid-price at time $t$ and the expectation is taken over the changes occuring due to trades alone) against varying relative tick sizes in Figure 3. On a log-log scale, again, we see a clear linear relationship between the two for small and medium tick stocks. There is some non-linearity for large tick stocks since the absolute returns are floored at half a cent. We show this non-linearity by depicting two regressions - one for large tick stocks (purple line) and one for all other stocks (red line). The slope of the fit for small/medium tick stocks is $-0.95$. We therefore state the following relation approximately:
$<r_{mid}>\propto\epsilon^{-1}$ (5)
Since relative tick-size here is proportional to the price of the security, we therefore conclude that the mean trade induced price is almost linearly related to the current price of the security. This experiment verifies the utility of using geometric scaling for price process models (the most simple one being the Geometric Brownian Motion).